OpenStudy (anonymous):

Let f(x) = |x+3 | / x+3 find each limit if possible lim x --> -3

myininaya (myininaya):

well the definition of |x+3|=x+3 if x+3>=0 =-(x+3) if x+3<0 or let's write it as |x+3|=x+3 if x>=-3 =-(x+3) if x<-3 Now since you have x approaching -3 and your function doesn't exist at x equal 3 you are going to need to look at the left and the right of -3 since we actually do have two different functions the function before -3 is -(x+3) and the function after -3 is x+3

OpenStudy (anonymous):

So you don't apply |u| = -u if u <0 ? It is the same function and I need to find the answers from the right and from the left, but all I get is undefined because 0/0

myininaya (myininaya):

I did apply |u|=-u if u<0 above

myininaya (myininaya):

you need to look at the right and left limit

OpenStudy (anonymous):

Exactly, so from the left it would be undefined? I'm having trouble understanding this chapter

myininaya (myininaya):

\[\lim_{x \rightarrow -3^+}\frac{|x+3|}{x+3} \text{ and } \lim_{x \rightarrow -3^-}\frac{|x+3|}{x+3}\] Now what is |x+3| for numbers to the right of -3? I actually already told you when I wrote the definition of |x+3|

myininaya (myininaya):